OpenStudy (anonymous):

Solve x(t) for 2x”(t)+6x’(t)+4x(t)=12, x(0)=3, x’(0)=1.

OpenStudy (jamesj):

You are going to have to learn some technique to be able to solve this. What have you been taught so far?

OpenStudy (jamesj):

Still there? Can't help if you don't talk to me.

OpenStudy (anonymous):

how do u rewrite the expression for 2x''(t)

OpenStudy (jamesj):

That's not the approach. What technique or techniques have you been taught for equations like this?

OpenStudy (anonymous):

well, i need to write the eq into standard form 6x'(t)= 6dx/dt what is 2x''(t)????

OpenStudy (jamesj):

x'' is the second derivative of the function x(t) with respect to t.

OpenStudy (jamesj):

In short the approach is this - solve the homogeneous equation 2x'' + 6x' + 4x = 0 * To do this, substitute x(t) = e^(rt) and find a quadratic equation in r * solve the equation to find values of r, call them r1 and r2 * then the general solution of the homogeneous equation is \[ x_h(t) = c_1 e^{r_1t} + c_2 e^{r_2t} \] - now find a particular solution to the inhomogeneous equation 2x'' + 6x' + 4x = 12 * I'll tell you what it is: \( x_p(t) = 3 \). Substitute it in and you'll see it works. - the general solution of the inhomogeneous equation is the sum of the homogeneous and particular solutions: \[ x(t) = x_h(t) + x_p(t) = c_1 e^{r_1t} + c_2 e^{r_2t} + 3 \] - finally, apply the initial conditions of \[ x(0)=3, \ \ x’(0)=1 \] to that general solution to solve for the constants \( c_1 \) and \( c_2 \) .